GLOBAL FREQUENCY AND INFRARED INTENSITY ANALYSIS OF ^{12}CH_{4} LINES IN THE 900–4800 CM^{−1} REGION

A. NIKITIN, Laboratory of Theoretical Spectroscopy, Institute of Atmospheric Optics, Russian Academy of Sciences, 634055 Tomsk, Russia; \underline{V. BOUDON}, J.-P. CHAMPION, LPUB – CNRS UMR 5027, 9 Av. A. Savary, BP 47870, F-21078 Dijon Cedex, France; S. ALBERT, S. BAUERECKER, M. QUACK, Physical Chemistry, ETH Zürich, CH-8093 Zürich, Switzerland; L. R. BROWN, Jet Propulsion Laboratory, California Institute of Technology, 4800 Oak Grove Drive, Pasadena, California 91109, USA.

A new global analysis of methane lines in the 900–4800 cm^{−1} region has been performed thanks to new experimental data for both line positions and intensities. This implies three of the ^{12}CH_{4} polyads, namely the dyad (940–1850 cm^{−1}, 2 vibrational levels, 2 sublevels), the pentad (2150–3350 cm^{−1}, 5 vibrational levels, 9 sublevels) and the octad (3550–4800 cm^{−1}, 8 vibrational levels, 24 sublevels) and some of the associated hot bands. New FTIR spectra of the pentad and octad regions have been recorded with a very high resolution (better than 0.001 cm^{−1} instrumental bandwidth, unapodized) at 78 K using the Bruker IFS 125 HR Zürich prototype (ZP2001) spectrometer^{a}. New intensity measurements were performed in the whole region at the Kitt Peak National Observatory. We also used previously recorded high-resolution Raman spectra^{b}. The effective Hamiltonian was expanded up to order 6 for the ground state, 6 for the dyad, 5 for the pentad and 5 for the octad. We obtain global root mean square deviations d_{RMS} for line positions = 1.4 × 10^{−4} cm^{−1} for the dyad, 6.0×10^{−4} cm^{−1} for the pentad and 3.3×10^{−3} cm^{−1} for the octad. This analysis represents a large improvement over the previous one^{c} with d_{RMS} = 0.041 cm^{−1} for the octad system.

aS. Albert ,K. Albert and M. Quack, TOPS (Trends in Optics and Photonics Series) 84, 177–180 (2003).

bD. Bermejo, J. Santos and P. Cancio, J. Mol. Spectrosc. 156 15–21 (1992) et J.-M. Jouvard, PhD Thesis, Dijon (1991).

cJ.-C. Hilico, O. Robert, M. Loëte, S. Toumi, A. S. Pine and L. R. Brown, J. Mol. Spectrosc. 208, 1–13 (2001).